OpenStudy (anonymous):

Help with Algebraic Symbol Manipulation: Please solve for x. T=1/6(x-y+z) x =

OpenStudy (anonymous):

x=6T+y-z

OpenStudy (anonymous):

Fantastic you were right! Thank you Ynooran! Tuiti Can you solve this on? Help with Algebraic Symbol Manipulation: Please solve for m. T=M/v2/r m =

OpenStudy (anonymous):

@Loser66 you missed the 6

OpenStudy (anonymous):

can i get medal?

OpenStudy (loser66):

how can?? don't understand. ynooran.

OpenStudy (anonymous):

m=T(v^2)/r

OpenStudy (loser66):

@ynooran how can you get x = 6T...

OpenStudy (anonymous):

ok: you divide both sides by 1/6

OpenStudy (anonymous):

that gives you 6T=(x-y+z)

OpenStudy (anonymous):

no wait I was right

OpenStudy (loser66):

oh, it is \[T = \frac{1}{6}*(x-y+z)\] I thought that \[T = \frac{1}{6(x-y+z)}\]

OpenStudy (loser66):

ok, you are right,!!! me stupid

OpenStudy (anonymous):

no youre smart you just missed something

OpenStudy (loser66):

thanks